OpenStudy (anonymous):

Solve x2 + 2x + 9 = 0 Solve x2 - 7x = -13 Solve -3x2 - 4x - 4 = 0 Solve 3x2 = -12x - 15 Solve -2x2 - 16x - 44 = 0

OpenStudy (anonymous):

Hint: Use the quadratic equation, it will solve for all these. Remember that a = the coefficient of the squared term (1 for the first one), b = the coefficient of the non-squared variable (2 for the first one) and c is the constant (9 for the first one). Then:\[x_{1/2} = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]Solves these quadratic.

OpenStudy (anonymous):

i already have all the soluations, i neeed my work checked.. and to compare anwsers

OpenStudy (anonymous):

bmp is right, you should use quadratic equation for this.

OpenStudy (anonymous):

You can write your steps, and we can see if they're correct.

OpenStudy (precal):

http://www.wolframalpha.com/input/?i=solve+x%5E2+%2B2x%2B9%3D0 use wolfram to check yourself